OpenStudy (anonymous):

What is the slope of a line defined by the equation 3x + y = -4?

Nnesha (nnesha):

y=mx+b is slope intercept form where m is slope and b is y-intercept you have to write 3x+y=-4 in slope intercept so solve for y

Nnesha (nnesha):

3x+y=-4 do you know how to solve for y ?

OpenStudy (anonymous):

No.

OpenStudy (anonymous):

-1?

Nnesha (nnesha):

no

Nnesha (nnesha):

\[\huge\rm 3x+y=-4\] move `3x` to the right side

OpenStudy (anonymous):

-7?

Nnesha (nnesha):

hmm how ?

Nnesha (nnesha):

how would you cancel out the 3x from the left side ? would u add 3x both sides ? or subtract 3x both sides ?